# Sensitivity of CTAO to axion-like particles from blazars: a machine learning approach

### v1

@@ Appendix B Classifier accuracy over the ALP parameter space
Appendix C Goodness of fit for the Beta distribution
@@
The example distributions of the $\Pi$ statistic shown in Figure 5 refer to cases where the accuracy of the classifiers is either worst or average. In these situations, the $\Pi$ values are more or less spread across the [0,1] interval and the resulting distribution is typically well fit by the Beta distribution of eq. (4.2). This can be checked e.g. by applying the Kolmogorov-Smirnov (KS) test [^12] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.kstest.html, which returns a large $p$-value when the data are compatible with the proposed distribution (i.e., the hypothesis that the data are described by said distribution is not rejected).
Figure 9: Example of a highly skewed $\Pi$ distribution obtained for a ML classifier close to maximal accuracy.
@@
In cases where the classifier accuracy is maximal, i.e. for large $g_{a\gamma}$ values, the resulting $\Pi$ distribution is largely peaked around very small values, as discussed in Section 4 (see e.g. Figure 9). In such extreme occurrences, the fit to the Beta distribution is typically not as good (resulting in small $p$-values from the KS test), along with being very sensitive to the numerical precision of the data. Moreover, the average value $\Pi_{0}$ obtained from the ALP-less datasets is usually close to 1, resulting in a confidence level for excluding the ALP hypothesis of essentially 100%, and consequently in a divergent $z$-score; see eqs. (3.8) and (3.9).
@@
In order to determine whether our main results hold under these limitations, we recomputed the exclusion confidence levels for each point in the ALP parameter space simply as the fraction of the 2000 simulated $\Pi(m_{a},g_{a\gamma})$ values smaller than $\Pi_{0}(m_{a},g_{a\gamma})$. This approach is in fact equivalent to eq. (3.8), but it does not require to assume a specific distribution for $\Pi$. By doing so, we have explicitly checked that the 95% CL ($2\sigma$) sensitivity region on the ALP parameter space remain unchanged. On the other hand, without the smoothing provided by a fitting function, the $z$-score diverges to infinity much more quickly as $g_{a\gamma}$ increases. We conclude that higher-confidence limits (e.g. $3\sigma$, $5\sigma$) obtained with the classifier-based method should be regarded as less reliable and more dependent on the quality of the fit.
Appendix D Discussion of systematic effects
Figure 10a shows the average $2\sigma$ sensitivity on the ALP parameter space from a 50-hour observation of PKS 2155-304 in its baseline state for different choices of the energy binning. We compare the results presented in the main text, obtained for 20 bins per decade, with those resulting from considering either half or double the number of bins per decade (10 or 40, respectively). The latter choice of binning, in particular, was employed in ref. [abdalla2021]. While there is very little difference between these results, halving the number of bins seems to reduce the size of the hole-like features around $g_{a\gamma}\sim 4\times 10^{-11}\,{\rm GeV}^{-1}$, while doubling it turns the feature into a continuous band. A similar outcome is found by comparing the results from different EBL models [franceschini2008, finke2010, gilmore2012] to the one used in the main text [dominguez2011], with the sensitivity to ALP parameters almost unaffected by this choice (Figure 10b).
Figure 10: CTAO sensitivity to the ALP parameter space obtained from a 50-hour simulated observation of PKS 2155-304 in its baseline state for different choices of the energy binning, EBL model, magnetic field strength and training strategy.

### v2

@@ Appendix C Classifier accuracy over the ALP parameter space
Appendix D Goodness of fit for the Beta distribution
@@
The example distributions of the $\Pi$ statistic shown in Figure 5 refer to cases where the accuracy of the classifiers is either worst or average. In these situations, the $\Pi$ values are more or less spread across the [0,1] interval and the resulting distribution is typically well fit by the Beta distribution of eq. (20). This can be checked e.g. by applying the Kolmogorov-Smirnov (KS) test [^12] https://docs.scipy.org/doc/scipy/reference/generated/
scipy.stats.kstest.html
, which returns a large $p$-value when the data are compatible with the proposed distribution (i.e., the hypothesis that the data are described by said distribution is not rejected).
@@
In cases where the classifier accuracy is maximal, i.e. for large $g_{a\gamma}$ values, the resulting $\Pi$ distribution is largely peaked around very small values, as discussed in Section 4 (see e.g. Figure 10). In such extreme occurrences, the fit to the Beta distribution is typically not as good (resulting in small $p$-values from the KS test), along with being very sensitive to the numerical precision of the data. Moreover, the average value $\Pi_{0}$ obtained from the ALP-less datasets is usually close to 1, resulting in a confidence level for excluding the ALP hypothesis of essentially 100%, and consequently in a divergent $z$-score; see eqs. (16) and (17).
@@
In order to determine whether our main results hold under these limitations, we recomputed the exclusion confidence levels for each point in the ALP parameter space simply as the fraction of the 2000 simulated $\Pi(m_{a},g_{a\gamma})$ values smaller than $\Pi_{0}(m_{a},g_{a\gamma})$. This approach is in fact equivalent to eq. (16), but it does not require to assume a specific distribution for $\Pi$. By doing so, we have explicitly checked that the 95% CL ($2\sigma$) sensitivity region on the ALP parameter space remain unchanged. On the other hand, without the smoothing provided by a fitting function, the $z$-score diverges to infinity much more quickly as $g_{a\gamma}$ increases. We conclude that higher-confidence limits (e.g. $3\sigma$, $5\sigma$) obtained with the classifier-based method should be regarded as less reliable and more dependent on the quality of the fit.
Figure 10: Example of a highly skewed $\Pi$ distribution obtained for a ML classifier close to maximal accuracy